The time spent watching TV exceeded the time spent listening to the radio by 581 hours. The combined time devoted to these two media was 2,529 hours. In 2007, how many hours did Americans spend listening to the radio and how many hours did they spend watching TV?
Found 2 solutions by mananth, stanbon:
Answer by mananth(16946)

let radio hours be x
TV hours = x+581
..
x+x+581=2529
2x+581=2529
add -581
2x+581-581=2529-581
2x=1948
divide by2
x= 1948/2
x= 974 hours - radio hours
TV hours = 2529-974 = 1555 hours

Answer by stanbon(75887)

The time spent watching TV exceeded the time spent listening to the radio by 581 hours.
Equations:
V = R+581
--------
The combined time devoted to these two media was 2,529 hours.
V + R = 2529
-------
In 2007, how many hours did Americans spend listening to the radio and how many hours did they spend watching TV?
---
Substitute for "V" and solve for "R":
R+581+R = 2529
2R = 1948
R = 974 hrs (time spent listening to radio)
---
Since V+R = 2529, V = 2529-974 = 1555 hrs (time spent on TV)
====================================
Cheers,
Stan H.
=============